40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 17 16:44:04 2015, Koji, Update, LSC, Delay line installed again 
    Reply  Tue Feb 17 19:49:51 2015, Koji, Update, LSC, Delay line un-installed again 
       Reply  Mon Feb 23 21:57:13 2015, Koji, Update, LSC, Delay line installed again (experiment, round 1) 
          Reply  Tue Feb 24 01:57:01 2015, Jenne, Update, LSC, Modelled effect of relative modulation phase ChangeInAngle_PRCLvsMICH.pngREFL33_MICH_SensMat_vsModulationOrder_vsModulationPhase.pngREFL33_SensMat_vsModulationPhase.pngREFL33_PRCL_nominalVsLoweredGamma.pngREFL33_MICH_nominalVsLoweredGamma.png
             Reply  Thu Feb 26 00:20:54 2015, ericq, Update, LSC, Modelled effect of relative modulation phase 33budget_canceled.png33budget_nominal.png2015-02-ModPhase.zip
                Reply  Thu Feb 26 01:51:39 2015, ericq, Update, LSC, Sideband HOMs C1_HOMcurves_PR.pngC1_HOMcurves_Y.pngC1_HOMcurves_X.pngC1_HOMlist.zip
                Reply  Thu Feb 26 01:53:35 2015, Koji, Update, LSC, Modelled effect of relative modulation phase 
       Reply  Tue Feb 24 23:39:16 2015, Jenne, Update, LSC, Delay line re-installed again 
          Reply  Wed Feb 25 04:59:47 2015, Jenne, Update, LSC, Delay line re-installed, measurements round 2 SensMat_PRMI_24Feb2015.pdf
             Reply  Wed Feb 25 12:16:27 2015, Koji, Update, LSC, Delay line re-installed, measurements round 2 
                Reply  Wed Feb 25 14:18:28 2015, rana, Update, LSC, Delay line re-installed, measurements round 2 
          Reply  Wed Feb 25 14:51:13 2015, Jenne, Update, LSC, Delay line un-installed again 
             Reply  Wed Feb 25 20:00:39 2015, Jenne, Update, LSC, Delay line un-installed again - sensing matrix comparisons 
                Reply  Wed Feb 25 23:48:57 2015, rana, Update, LSC, some thoughts 
                   Reply  Thu Feb 26 20:02:46 2015, Jenne, Update, LSC, re: some thoughts 
Message ID: 11059     Entry time: Mon Feb 23 21:57:13 2015     In reply to: 11045     Reply to this: 11060
Author: Koji 
Type: Update 
Category: LSC 
Subject: Delay line installed again (experiment, round 1) 

Last Wednesday we tried PRMI 3f modulation cancellation. Under the 3f modulation cancellation, the PRMI could not be locked
with REFL signals, and the PRCL signal was just barely sufficient to lock PRCL with help of AS55Q MICH.

- The PRCL signal level in REFL33 was reduced by factor of 20 compared with the conventional modulation setting.
=> The 3f modulation cancellation does not chage the level of 11/22MHz sidebands, it is expected that REFL33 signal
has no significant change of the signal level. But it does.  If we change the relative phase between the modulations
at 11 and 55MHz, the signal level is recovered by factor of 5. Therefore something related to 55MHz modulation
(55MHz x 22MHz, or 44MHz x 11MHz) was contributing more than -11MHzx22MHz.

- Under the 3f demodulation cancellation, MICH signal in the REFL ports were extremely weak and there was
no hope to use it for any feedback control.

- WIth the PRMI locking by REFL33->PRCL and AS55Q->MICH, the sensing matrix was measured. All of the REFL
ports however, showed extremely degenerate sensing matrix between MICH and PRCL.

This was enough confusing to us, and we didn't draw any useful information from these. Here are some ideas to
investigate what is happening in out optical and electrical system.

- One approach is to use as simple optical setup as possible to inspect our sensing systems. For example,
we may want to try PRX/PRY/XARM/YARM cavities to check the functions of the REFL diodes and qualitatively characterize
the sensing chain (Optical gain [W/m], noise level, demodulation phase) so that we can compare these with
an interferometer seinsing model.

- Another approach is to change the mdulation setting more freely and empirically try to find the characteristic
of our optical/electrical systems. e.g. change the relative modulation phase and/or 55MHz attenuation, and try to understand
how 11-22, 11-44, 22-55, 0-33 pairs are contributing the signal.

ELOG V3.1.3-